OpenStudy (anonymous):

y' = x-2y cot 2x ,,,, using linear differential equation..? help?

OpenStudy (unklerhaukus):

\[y' = x-2y \cot 2x\]\[y' +2y \cot 2x= x\]\[\mu(x)=e^{\int \cot 2x\text dx}\]

OpenStudy (anonymous):

go on..

OpenStudy (anonymous):

so whats the right answer?

OpenStudy (unklerhaukus):

hane you found the integrating factor \(\mu(x)\) ?

OpenStudy (anonymous):

no,,have you found it ?

OpenStudy (unklerhaukus):

yes

OpenStudy (anonymous):

shoe it please..

OpenStudy (anonymous):

show it please

OpenStudy (unklerhaukus):

\[\mu(x)=e^{\int \cot 2x\text dx}=\dots\]

OpenStudy (anonymous):

then what is the correct anwer?

OpenStudy (unklerhaukus):

what step are you having difficulty on

OpenStudy (anonymous):

in finding the General Formula ..

OpenStudy (unklerhaukus):

\[\int \cot 2x \text dx=\frac 12\int \cot u\text du=\cdots\]

OpenStudy (unklerhaukus):

that's not a step,

OpenStudy (anonymous):

so whats the general formula ?

OpenStudy (anonymous):

please give me the right answer? its kinda hard to solve for me

OpenStudy (unklerhaukus):

im not going to solve the entire problem for you, but i can help you with specific steps , if you'd rather help from a robot, try this http://www.wolframalpha.com/input/?i=y%27+%3D+x-2y+cot+2x

OpenStudy (anonymous):

ok thanks for helping